OpenStudy (anonymous):

Check my answer! Medal! my answer is x+3/(x-6)(x+7) question below

OpenStudy (anonymous):

add or subtact

OpenStudy (anonymous):

@amistre64 would you mind checking this please?

OpenStudy (amistre64):

the denom is a good approach, lets check the top does: x(x+7) + 3(x-6) = x+3? x^2+7x + 3x-18 = x+3 x^2+10x-18 = x+3 ... no, it does not

OpenStudy (anonymous):

ive tried wolfalpha but it doesnt give me the answer I am looking for @dumbcow and ok what do I do next then @amistre64

OpenStudy (amistre64):

well, when x=0, -18 not= 3 your solution is not correct ... correct the solution, and then you can see if the top can factor to reduce some stuff top to bottom

OpenStudy (anonymous):

ok one sec I am gonna try it again

OpenStudy (anonymous):

ok now I got x^2+8x-18/(x-6)(x+7)

OpenStudy (dumbcow):

i dont know what you are looking for but it certainly does provide the answer @amistre64 has also shown you the correct way to get the answer

OpenStudy (anonymous):

I agree it provides the correct answer but my answers for some reason are listed differently and I don't have that option in my answers

OpenStudy (anonymous):

@amistre64

OpenStudy (dumbcow):

then there is a problem with the given answers or they are just written or simplified differently what answers do they provide?

OpenStudy (dumbcow):

In general: \[\frac{A}{x-6} + \frac{B}{x+7} = \frac{A(x+7) + B(x-6)}{(x-6)(x+7)}\]

OpenStudy (dumbcow):

umm the answer provided by wolfram and amistre64 is listed ... please check again

OpenStudy (anonymous):

so the one on wolfram is just simplified?

OpenStudy (anonymous):

oh i see it haha nvm thanks